OpenStudy (anonymous):

trying to solve 3x/5 - 2 = x + 3

OpenStudy (anonymous):

For starters, you could move the 2 over to the right. Do you know how to do that?

OpenStudy (anonymous):

I do but I just went another route. 15 is the common denominator of 5 and 3 so I did the following. What do you think,,, 15(3x/5 - 2) = 15(x + 1/3) 15 . 3x/5 - 15(2) = 15 (x) + 15 (1/3) 9x - 30 = 15x + 5 9x - 30 - 9x = 15x + 5 - 9x -30 - 5 = 6x + 5 - 5 -35 = 6x x= -35/6 Is that right?

OpenStudy (anonymous):

That would be correct. However, note that in your original wording you had (x + 3) on the right side but in your last post you put (x + 1/3). If it really is 1/3, then your answer is correct. If you made a typo and it's supposed to be 3 then your answer would not be correct but your method is right.